# Latent Multimodal Reconstruction for Misinformation Detection

### v1

@@ Abstract
Multimodal misinformation, such as miscaptioned images, where captions misrepresent an image’s origin, context, or meaning, poses a growing challenge in the digital age. To support fact-checkers, researchers have been focusing on creating datasets and developing methods for multimodal misinformation detection (MMD). Due to the scarcity of large-scale annotated MMD datasets, recent studies leverage synthetic training data via out-of-context image-caption pairs or named entity manipulations; altering names, dates, and locations. However, these approaches often produce simplistic misinformation that fails to reflect real-world complexity, limiting the robustness of detection models trained on them. Meanwhile, despite recent advancements, Large Vision-Language Models (LVLMs) remain underutilized for generating diverse, realistic synthetic training data for MMD. To address this gap, we introduce “MisCaption This!”, a training dataset comprising LVLM-generated miscaptioned images. Additionally, we introduce “Latent Multimodal Reconstruction” (LAMAR), a network trained to reconstruct the embeddings of truthful captions, providing a strong auxiliary signal to the detection process. To optimize LAMAR, we explore different training strategies (end-to-end training and large-scale pre-training) and integration approaches (direct, mask, gate, and attention). Extensive experiments show that models trained on “MisCaption This!” generalize better on real-world misinformation, while LAMAR sets new state-of-the-art on both NewsCLIPpings and VERITE benchmarks; highlighting the potential of LVLM-generated data and reconstruction-based approaches for advancing MMD. We release our code at: https://github.com/stevejpapad/miscaptioned-image-reconstruction.
@@ 1 Introduction
The rise of the internet and digital technologies has significantly accelerated information dissemination while also amplifying the spread of misinformation, enabling new forms of deceptive content such as DeepFakes [1], multimodal misinformation [2], and LLM-generated misinformation [3]. Given the scale and rapid spread of misinformation, researchers have been developing automated fact-checking methods to support human fact-checkers in identifying false information more efficiently [4]. In this study, we focus on multimodal misinformation detection (MMD), specifically, the detection of misleading image-caption pairs, where texts and images jointly contribute to the spread of false or deceptive information [5].
Recent studies on MMD primarily focus on developing large-scale datasets and detection methods. Aside from a few small-scale annotated datasets [6, 7], existing datasets primarily consist of weakly annotated [8, 9, 10] or algorithmically generated datasets; created either by pairing images with out-of-context (OOC) captions from other images [11, 12, 13, 14] or by manipulating named entities to introduce inconsistencies, resulting in miscaptioned (MC) images [15, 16, 17]. On the modeling side, detection models have leveraged large pre-trained backbone encoders [14] or fine-tuned through self-supervised learning [18], improved modality fusion with attention-based techniques [19, 20, 21], incorporated external evidence [22, 23], and utilized Large Vision-Language Models (LVLMs) for both detection [24, 25] and explanation generation [26].
Nevertheless, named-entity manipulations often produce simplistic misinformation, lacking basic factual or logical consistency, while the potential of LVLMs to generate more diverse and robust synthetic training data remains unexplored. Moreover, despite the success of reconstruction networks in other domains, their potential remains largely unexplored for MMD where is particularly relevant, as the task closely aligns with human fact-checking practices, where reconstructing the true origin, context, or meaning of an image is a common strategy for debunking misinformation [27].
@@
Figure 1: High-level overview of the proposed framework: an LVLM modifies a truthful image-caption pair, while LAMAR’s reconstruction network re-creates the original caption embedding produced by CLIP. This embedding is fused with other modalities using a mechanism (e.g., Gate) and fed into the detection network to produce the final verdict. The reconstruction network is trained to minimize the error between original and reconstructed embeddings, while the detection network is optimized for classification.
To this end, we propose a framework that leverages an LVLM to manipulate the original captions of images sourced from a dataset of truthful image-caption pairs, generating false captions that misrepresent aspects of the images. Subsequently, the proposed Latent Multimodal Reconstruction (LAMAR) network is tasked with reconstructing the embeddings of the original, truthful captions. As shown in Fig. 1, a truthful image-caption pair is manipulated by the LVLM, while LAMAR is tasked with reconstructing the original caption embedding using the image and the generated caption embeddings. Thereafter, the reconstructed embedding is integrated into the final detection network alongside the fused modalities. Our rationale is that LVLMs can generate more diverse and realistic training data, enhancing the generalizability of detection models, while the reconstruction process serves as an auxiliary signal, leveraging learned representations to refine embeddings by capturing consistency and inconsistency patterns in image-caption pairs—even in the absence of external evidence.
@@
More specifically, we leverage an LVLM alongside “adversarial prompt selection” to limit the number of generative prompts by evaluating them against the LVLM’s zero-shot detection capabilities; filtering out prompts that produce easily detectable misinformation or overly generic image descriptions. This process results in the selection of four generative prompts, each used to generate a distinct version of our “MisCaption This!” dataset. For LAMAR, we employ a Transformer encoder with element-wise vector operations for modality fusion and explore two training strategies, end-to-end training and large-scale pre-training, as well as four methods for integrating the reconstructed embeddings into the detection network: direct integration, masking, gating, and self-attention.
Through extensive experiments with various training datasets, we demonstrate that models trained on “MisCaption This!” achieve superior out-of-distribution generalization on the VERITE evaluation benchmark [28], outperforming those trained on datasets relying on entity manipulation or cross-modal misalignment—with LAMAR performing 7.8% and 10.4% better, respectively. Moreover, LAMAR consistently outperforms prior state-of-the-art (SotA) models across training settings and datasets, achieving improvements of 4.3% on ‘True vs. MC’, 3.0% on ‘True vs. OOC’, and 4.8% on the multiclass task when trained on “MisCaption This!”.
@@ 2 Related Work
In recent years, automated fact-checking has attracted growing research interest [4], encompassing a range of challenging tasks such as claim detection [29], evidence filtering and retrieval [30, 31], fake news detection [32, 33, 34], retrieval of fact-checked articles [35, 36], DeepFake detection [1], and multimodal forms of misinformation [2]. In this paper, we focus on MMD, which is garnering increasing attention from researchers, leading to the ongoing development of datasets and methodologies for detecting false information and inconsistencies between images and their accompanying textual captions.
@@ 2.1 MMD Datasets
Training machine learning models for MMD requires suitable datasets, with current research focusing on annotated, weakly annotated, and synthetically generated data. Early MMD datasets, such as the ‘Twitter’ [7] and ‘Weibo’ [37] datasets, are relatively small and cover only a limited number of events-17 and 73, respectively-raising concerns about model generalization. To address this, larger weakly annotated datasets have emerged, including MuMiN [9], with rich social context but few images, and NewsBag [10], which includes satirical content, and Fakeddit [8], with over a million instances collected from Reddit. However, studies show that models trained on ‘Twitter’ and Fakeddit often exhibit unimodal biases, undermining their effectiveness in real-world multimodal misinformation detection [28].
@@
NES-based datasets generate misinformation by replacing named entities in captions with alternatives retrieved from similar or contextually relevant texts using cluster-based retrieval (MEIR [15]), rule-based substitutions (TamperedNews [16]), and CLIP-based retrieval (CLIP-NESt [17]). Despite advancements in LVLMs, only MMFakeBench leverages LVLM-generated rumors and AI-altered images and serves as a small evaluation benchmark [38]. Instead, we introduce a large LVLM-generated training dataset.
@@ 2.2 MMD Methods
Research on MMD has centered on developing models that encode textual and visual modalities, fuse their representations, and assess their consistency and factual accuracy. Early methods, such as SpotFake [39], used VGG-19 and BERT, while recent approaches use pre-trained multimodal encoders such as CLIP [14], or fine-tune it through self-supervised learning [18]. Some models incorporate multi-task learning, such as EANN with an event discriminator [40] or MVAE, which uses an autoencoder to reconstruct the input text and visual features [41]; but does not modify the input text or reconstruct truthful captions from false ones.
While earlier methods relied on simple concatenation of visual and textual embeddings, more advanced approaches have explored Attention-based Multimodal Bilinear Pooling [19], Bidirectional Crossmodal Fusion (BCMF) [20], multi-head attention in Transformers [17], and element-wise vector fusion [21] to enhance cross-modal interaction. Recent work integrates external web evidence, with methods assessing internal and external consistency (CCN [22], SNIFFER [24]) or evaluating stance and relevance of external evidence (SEN [23], RED-DOT [21]). While external evidence is shown to improve performance, concerns remain about ‘leaked evidence’ from fact-checking articles [42, 31] and dataset artifacts that models may exploit instead of assessing factuality [43]. For these reasons, we do not consider evidence-based approaches in this study.
@@ 2.3 Reconstruction Networks
Reconstruction networks are deep learning models designed to generate or restore original data from low-resolution or altered inputs. They have been applied in various domains, including few-shot image classification by reframing it as a reconstruction problem in latent space [44], image inpainting through an adversarial framework guided by textual descriptions [45], super-resolution reconstruction to enhance low-resolution images of the same scene [46], and DeepFake detection [47]. Reconstruction networks remain largely unexplored in MMD, with MVAE [41] being a notable exception. However, MVAE uses an autoencoder to reconstruct the input text and image embeddings from a joint latent space but does not modify the text or generate truthful image descriptions from false ones; its objective is to extract more informative features for MMD within the latent space.
@@ 3 Problem Formulation
Given a set $\mathcal{D}^{t}={(I^{t}_{i},C^{t}_{i})}_{i=1}^{N}$ of $N$ image-caption pairs, where $I^{t}_{i}$ and $C^{t}_{i}$ represent an image and its matching, truthful caption, we define ‘Out-Of-Context’ (OOC) as any pair that combines an image $I^{t}_{i}$ with a caption $C^{x}_{i}$ taken from another image within $\mathcal{D}^{t}$; where x denotes the context. Similarly, we define ‘Mis-Captioned’ (MC) as any image whose accompanying caption $C^{f}_{i}$ has been manipulated to misrepresent the original, content, and/or meaning of the image; where “f” denotes falsehood.
We define a “Manipulator” as any method used to generate OOC pairs (e.g., CLIP-based retrieval) or MC pairs (e.g., named entities manipulation) from the original set $\mathcal{D}^{t}$ set, resulting in the sets $\mathcal{D}^{x}$ and $\mathcal{D}^{f}$, respectively.
We define MMD as a classification task with the objective to learn a mapping function $\mathsf{M}^{d}:\mathcal{D}\rightarrow\hat{y}$ where $\hat{y}$ represents the prediction upon target class $y$ in either of three scenarios:
1) Binary: ‘True vs. MC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{f}]$ with $K=N*2$ total pairs and $y\in\{0,1\}$
2) Binary: ‘True vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x}]$ with $K=N*2$ total pairs and $y\in\{0,2\}$,
3) Multi-class: ‘True vs. MC vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x},\mathcal{D}^{f}]$ with $K=N*3$ total pairs and $y\in\{0,1,2\}$.
Given image and text encoders $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$ producing image embeddings $\mathbf{I}$ and text embeddings $\mathbf{C}$ for image-caption pairs $(I,C)$ under examination, we define latent reconstruction as the task of learning a mapping function $\mathsf{M}^{r}:(\mathbf{I},\mathbf{C})\rightarrow\mathbf{\hat{C}^{t}}$, where $\mathbf{\hat{C}^{t}}$ represents the predicted embedding of the original, truthful caption $\mathbf{C^{t}}$. The model $\mathsf{M}^{r}$ is trained to minimize the reconstruction error between the true and predicted caption embeddings, such that:
@@
where $\mathcal{L}_{r}$ is the loss function (i.e., the Mean Squared Error (MSE)) that quantifies the discrepancy between the embeddings of the true and reconstructed caption embeddings.
Figure 2: Examples of truthful and generated captions taken from the “MisCaption This!” ($\mathcal{D}_{3}$) dataset, alongside false captions created via named entity swaps.
4 Construction of “MisCaption This!”
In this study, we explore the creation of a synthetic training dataset of miscaptioned images ($\mathcal{D}^{f}$) by manipulating the image captions of a truthful dataset ($\mathcal{D}^{t}$) using an LVLM as the “Manipulator”. Our rationale is that LVLMs, with their advanced multimodal understanding and generation capabilities, can produce more realistic false captions for images compared to methods relying on manipulating named entities. In turn, we hypothesize that the generated data $\mathcal{D}^{f}$ can be leveraged to train more robust detection models $\mathsf{M}^{d}$, thus enhancing their ability to generalize to real-world misinformation.
@@ 4.1 Generative Model
To generate synthetic data, we employ LLaVa-1.6 (Large Language and Vision Assistant) [48] from Hugging Face [^1] https://huggingface.co/llava-hf/llava-v1.6-mistral-7b-hf leveraging Mistral-7B-Instruct-v0.2. We also explored LLaVa-1.6 (Vicuna-13B), Janus Pro 7B (DeepSeek), and MiniGPT-v2 (Llama-2-7B) as alternatives, however, we were unable to get them to consistently generate realistic false captions, as they often defaulted to generic image captioning, re-phrasing of the original caption, or overly simplistic misinformation. In contrast, GPT-4o Mini demonstrated robust safeguards, often refraining from generating misinformation altogether. Therefore, this part of our work serves as an initial exploration of LVLM-generated synthetic data, with future research needed to assess the impact of other LVLMs and prompting methods.

### v3

# Latent Reconstruction from Generated Data for Multimodal Misinformation Detection

@@ Abstract
Multimodal misinformation, such as miscaptioned images, where captions misrepresent an image’s origin, context, or meaning, poses a growing challenge in the digital age. Due to the scarcity of large-scale annotated datasets for multimodal misinformation detection (MMD), recent approaches rely on synthetic training data created via out-of-context pairings or named entity manipulations (e.g., altering names, dates, or locations). However, these often yield simplistic, unrealistic examples, which limits their utility as training examples. To address this, we introduce “MisCaption This!”, a framework for generating high-fidelity synthetic miscaptioned datasets through Adversarial Prompting of Vision-Language Models (VLMs). Additionally, we introduce “Latent Multimodal Reconstruction” (LAMAR), a Transformer-based network trained to reconstruct the embeddings of truthful captions, providing a strong auxiliary signal to guide detection. We explore various training strategies (end-to-end vs. large-scale pre-training) and integration mechanisms (direct, mask, gate, and attention). Extensive experiments show that models trained on “MisCaption This!” data generalize better to real-world misinformation, while LAMAR achieves new state-of-the-art on NewsCLIPpings, VERITE, and the newly introduced VERITE 24/25 benchmark; highlighting the efficacy of VLM-generated data and reconstruction-based networks for advancing MMD. Our code is available at https://github.com/stevejpapad/miscaptioned-image-reconstruction.
@@ 1 Introduction
The rise of the Internet and digital technologies has significantly accelerated the spread of information but also the proliferation of misinformation, including new forms of deceptive content such as DeepFakes [32], multimodal misinformation [2], and LLM-generated misinformation [8]. Given the scale and speed of misinformation dissemination, researchers are developing automated fact-checking tools to assist human fact-checkers in identifying deceptive content more efficiently [11]. In this study, we focus on multimodal misinformation detection (MMD), specifically, the detection of misleading image-caption pairs, where both modalities jointly contribute to the spread of false or misleading information [3].
Due to the scarcity of large-scale, expert-annotated MMD datasets, researchers have increasingly turned to synthetic training data. While fact-checking archives provide high-fidelity labels, their limited volume of multimodal pairs, lack of negative samples, and licensing constraints render them useful for evaluation but insufficient for training. Consequently, aside from a few small-scale annotated resources [44], recent studies have had to rely on either weakly annotated [6, 25, 26, 15] or algorithmically generated training datasets; created either by pairing images with out-of-context (OOC) captions from other images [13, 4, 5, 22] or by manipulating named entities to introduce inconsistencies, resulting in miscaptioned (MC) images [33, 24, 30]. However, named-entity manipulation often produces shallow or implausible misinformation, lacking the nuance seen in real-world examples. At the same time, the potential of Vision Language Models (VLMs) to generate richer, more diverse synthetic training data remains underexplored.
On the modeling side, recent works have explored large pre-trained encoders [22], self-supervised fine-tuning [23], attention-based fusion mechanisms [18, 40, 27], external evidence [1, 41], and Vision-Language Models (VLMs) for both detection [31, 35] and explanation generation [42]. However, while reconstruction networks have shown promise in other domains, their use in MMD is still limited – despite their alignment with human fact-checking strategies, where reconstructing the original meaning, context, or source of an image is key to detecting misinformation [7].
@@
Figure 1: High-level overview of the proposed workflow. A VLM generates a false caption from the original, truthful image-caption pair. The Reconstructor (Transformer encoder) then takes both the image and false caption embeddings as inputs to identify and rectify inaccuracies in the text, recreating the original truthful caption embedding. This reconstructed representation is fused with the other modalities via a specialized mechanism (e.g., Gating or Attention) and passed to the detector to produce the final verdict. The network is trained to simultaneously minimize the error between original and reconstructed embeddings and optimized for classification accuracy.
To this end, we propose “MisCaption This!”, a framework that leverages VLMs to manipulate the captions of images drawn from truthful image-caption datasets, generating synthetic false captions that misrepresent aspects of the image. The resulting datasets are used to train the proposed Latent Multimodal Reconstruction (LAMAR) network, which learns to reconstruct the embedding of the original, truthful caption. As shown in Fig. 1, LAMAR takes the image and manipulated caption embeddings as input and is tasked with recovering the original caption embedding. This reconstructed embedding is then integrated into the final detection network alongside the fused multimodal representation.
Our rationale is that (1) VLMs can generate more diverse and realistic synthetic training data, and (2) semantic divergence between reconstructed and input embeddings provides a discriminative signal that exposes cross-modal and intra-modal inconsistencies, significantly improving model generalization.
@@
“MisCaption This!” uses open-source VLMs (LLaVA, Llama, Molmo) and apply adversarial prompt selection, where generative prompts are evaluated against the VLM’s zero-shot detection capabilities. Prompts that produce easily detectable or overly generic captions are discarded. LAMAR employs a Transformer encoder with element-wise modality fusion, and we explore both end-to-end training and large-scale pre-training. We also compare four integration strategies for the reconstructed embedding: direct integration, masking, gating, and self-attention.
Our main contributions can be summarized as follows:
  • We introduce “MisCaption This!”, a framework for creating synthetic training datasets of miscaptioned images using open-source VLMs.
  • We propose the Latent Multimodal Reconstruction (LAMAR) architecture, which learns to reconstruct the truthful caption embedding from its input.
  • We show that models trained on VLM-generated data generalize better to real-world misinformation, outperforming named-entity swapping and cross-modal misalignment datasets by 7.8% and 10.4% on the VERITE benchmark [28], respectively.
  • LAMAR achieves new state-of-the-art (SotA) results on NewsCLIPpings and across all VERITE tasks, with improvements of 4.3% on “True vs. MC”, 3.0% on “True vs. OOC”, and 5.6% on multiclass classification.
  • LAMAR demonstrates stronger temporal generalization and robustness on emerging misinformation compared to zero-shot VLMs, maintaining high performance (70.0%) on the new VERITE 24/25 benchmark – consisting of events after the VLM knowledge-cutoff – while VLM performance degrades significantly (8–12% reduction).
@@ 2.1 MMD Datasets
Training machine learning models for MMD requires suitable datasets, with current research focusing on annotated, weakly annotated, and synthetically generated data. Early MMD datasets, such as the Twitter [6] and Weibo [14] datasets, are relatively small and cover only a limited number of events-17 and 73, respectively-raising concerns about model generalization. To address this, larger weakly annotated datasets have emerged, including MuMiN [26], with rich social context but few images, and NewsBag [15], which includes satirical content, and Fakeddit [25], with over a million instances collected from Reddit. However, studies show that models trained on Twitter and Fakeddit often exhibit unimodal biases, undermining their effectiveness in real-world multimodal misinformation detection [28].
@@
NES-based datasets generate misinformation by replacing named entities in captions with alternatives retrieved from similar or contextually relevant texts using cluster-based retrieval (MEIR [33]), rule-based substitutions (TamperedNews [24]), and CLIP-based retrieval (CLIP-NESt [30]). Despite recent progress in VLMs and their use for synthetic dataset generation [19], their application to MMD remains limited. To date, only MMFakeBench [21] uses VLM-generated rumors and AI-manipulated images, and serves solely as a small-scale evaluation benchmark. In contrast, we introduce a large-scale VLM-generated training dataset for MMD.
@@ 2.2 MMD Methods
Research on MMD has centered on developing models that encode textual and visual modalities, fuse their representations, and assess their consistency and factual accuracy. Early methods, such as SpotFake [34], used VGG-19 and BERT, while recent approaches use pre-trained modality-specific dual-encoder models such as CLIP [22], or fine-tune it through self-supervised learning [23]. Some models incorporate multi-task learning, such as EANN with an event discriminator [36] or MVAE, which uses an autoencoder to reconstruct the input text and visual features [17]; but does not modify the input text or reconstruct truthful captions from false ones.
While earlier methods relied on simple concatenation of visual and textual embeddings, more advanced approaches have explored Attention-based Multimodal Bilinear Pooling [18], Bidirectional Crossmodal Fusion (BCMF) [40], multi-head attention in Transformers [30], and element-wise vector fusion [27] to enhance cross-modal interaction. Recent work integrates external web evidence, with methods assessing internal and external consistency (CCN [1], SNIFFER [31]) or evaluating stance and relevance of external evidence (SEN [41], RED-DOT [27]). While external evidence is shown to improve performance, concerns remain about ‘leaked evidence’ from fact-checking articles [10, 9] and datasets-specific artifacts that models may exploit instead of assessing factuality [29]. For these reasons, we do not consider evidence-based approaches in this study.
@@ 2.3 Reconstruction Networks
Reconstruction networks are deep learning models designed to generate or restore original data from low-resolution or altered inputs. They have been applied in various domains, including few-shot image classification by reframing it as a reconstruction problem in latent space [37], image inpainting through an adversarial framework guided by textual descriptions [38], super-resolution reconstruction to enhance low-resolution images of the same scene [39], and DeepFake detection [12]. Reconstruction networks remain largely unexplored in MMD, with MVAE [17] being a notable exception. However, MVAE employs a variational autoencoder to reconstruct the input, recovering the input words (via Bi-LSTMs) and VGG-19 image embeddings from a shared latent space. Crucially, this reconstruction serves only to enhance feature extraction and does not involve modifying or correcting the input content, e.g., generating truthful descriptions, or their corresponding latent representations, from false ones.
@@ 3 Problem Formulation
Given a set $\mathcal{D}^{t}={(I^{t}_{i},C^{t}_{i})}_{i=1}^{N}$ of $N$ image-caption pairs, where each $I^{t}_{i}$ is an image and $C^{t}_{i}$ its matching, truthful caption, we define:
Out-of-Context (OOC) pairs as $(I^{t}_{i},C^{x}_{i})$, where $C^{x}_{i}$ is a caption taken from a different sample in $\mathcal{D}^{t}$;
Mis-Captioned (MC) pairs as $(I^{t}_{i},C^{f}_{i})$, where $C^{f}_{i}$ is a manipulated version of the original caption that misrepresents the content or meaning of $I^{t}_{i}$.
A “manipulator” refers to any method used to create OOC (e.g., CLIP-based retrieval) or MC (e.g., entity swapping) samples, producing datasets $\mathcal{D}^{x}$ and $\mathcal{D}^{f}$ from $\mathcal{D}^{t}$.
We define MMD as a classification task to learn a mapping $\mathsf{M}^{d}:\mathcal{D}\rightarrow\hat{y}$, where $\hat{y}$ is the predicted class label under one of three settings:
(1) Binary: ‘True vs. MC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{f}]$ with K=N2K=N*2 total pairs and $y\in\{0,1\}$
(2) Binary: ‘True vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x}]$ with K=N2K=N*2 total pairs and $y\in\{0,2\}$,
(3) Multi-class: ‘True vs. MC vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x},\mathcal{D}^{f}]$ with K=N3K=N*3 total pairs and $y\in\{0,1,2\}$.
Let $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$ denote the image and text encoders, producing embeddings $\mathbf{I}$ and $\mathbf{C}$ for input pair $(I,C)$, respectively. We define latent reconstruction as learning a function $\mathsf{M}^{r}:(\mathbf{I},\mathbf{C})\rightarrow\hat{\mathbf{C}}^{t}$, where $\hat{\mathbf{C}}^{t}$ approximates the embedding of the original truthful caption $\mathbf{C}^{t}$. The model is trained to minimize reconstruction loss:
@@
where $\mathcal{L}_{r}$ is the Mean Squared Error (MSE) between the true and reconstructed caption embeddings.
Unless explicitly denoted as $C^{t}$ (truthful), $C^{f}$ (falsified), or $\mathbf{\hat{C}}^{t}$ (reconstructed), we use $C$ and $\mathbf{C}$ to refer to the input caption and its embedding–regardless of truthfulness.
4 “MisCaption This!”
In this study, we explore the creation of synthetic training datasets of miscaptioned images ($\mathcal{D}^{f}$) by manipulating the image captions of a truthful dataset ($\mathcal{D}^{t}$) using a VLM as the “Manipulator”. Our rationale is that VLMs, with their advanced multimodal understanding and generation capabilities, can produce more realistic false captions for images compared to methods relying on manipulating named entities. In turn, we hypothesize that the generated data $\mathcal{D}^{f}$ can be leveraged to train more robust detection models $\mathsf{M}^{d}$, thus enhancing generalization to real-world misinformation.
@@ 4.1 Generative Model
To generate synthetic training data, we employ open-source VLMs: LLaVA-1.6-Mistral [^1] https://huggingface.co/LLaVA-hf/LLaVA-v1.6-mistral-7b-hf, Molmo-7B-D [^2] https://huggingface.co/allenai/Molmo-7B-D-0924, and Llama-3.2-11B-Vision [^3] https://huggingface.co/meta-llama/Llama-3.2-11B-Vision. We also explored LLaVA-1.6-Vicuna-13B, Janus Pro 7B (DeepSeek), and MiniGPT-v2-Llama-2-7B as alternatives, however, we were unable to get them to consistently generate realistic false captions, as they often defaulted to generic image captioning, re-phrasing of the original caption, or overly simplistic misinformation. In contrast, GPT-4o Mini demonstrated robust safeguards, often refraining from generating misinformation altogether.